OpenStudy (buttermequeasy):

The formula for degrees Celsius is: C = 9 over 5 (F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the formula for F. Show all steps. Part 2: Determine how many degrees Fahrenheit 2 degrees Celsius is.

OpenStudy (anonymous):

\[C = \frac{9}{5}(F-32)\]\[\frac{5}{9}C = F - 32\]\[F = \frac{5}{9}C + 32\]

OpenStudy (anonymous):

I thought Celcius equaled \[C = \frac{5}{9}(F-32)\]

OpenStudy (anonymous):

Because \[F = \frac{9}{5}C + 32\]

OpenStudy (anonymous):

2 Celsius = 35.6 F